OpenStudy (anonymous):

Use the distance formmula to find the distance between point B (-3,-8) and point C (2,4).

OpenStudy (phoenixfire):

Do you know what the distance formula is?

OpenStudy (anonymous):

yes but I'm confused as to where to put the numbers in.

OpenStudy (phoenixfire):

\[d=\sqrt{x^2+y^2}=\sqrt{(x_1-x)^2+(y_1-y)^2}\] You have two points, find the change in x and change in y, and put into the formula.

OpenStudy (anonymous):

is -3 -8 x1 and y1

OpenStudy (phoenixfire):

it shouldn't make a difference since it's squared. But usually if you want to find the distance from B to C, your x1y1=C and xy=B

OpenStudy (anonymous):

Is the answer 17

OpenStudy (phoenixfire):

\[dist=\sqrt{(2+3)^2+(4+8)^2}=\sqrt{25+144}=\sqrt{169}=13\]

OpenStudy (anonymous):

My distance formula says you subtract x 2 and x 1

OpenStudy (phoenixfire):

yup. \[2-(-3)=2+3=5\]

OpenStudy (phoenixfire):

\[4-(-8)=4+8=12\] Double negative becomes a positive.

OpenStudy (anonymous):

Ok. Thanks for youre help.